OpenStudy (anonymous):

Solve for a. 48 6 - 7a A. a - 54 7 B. a -6 C. a 6 55 D. a 6

OpenStudy (anonymous):

Can you please put this in a readable format?

OpenStudy (anonymous):

A. a >or = too -54/7 B. a > or = too -6 C. a > or = too 6/55 D. a < or = too 6

OpenStudy (notamathgenius):

@brandbear All you do is take the numbers like A and plug it in and get the answer MAKE SURE YOU MULTIPLY FIRST